# Underwater Image Enhancement via Dehazing and Color Restoration

### v1

@@ Abstract
With the rapid development of marine engineering projects such as marine resource extraction and oceanic surveys, underwater visual imaging and analysis has become a critical technology. Unfortunately, due to the inevitable non-linear attenuation of light in underwater environments, underwater images and videos often suffer from low contrast, blurriness, and color degradation, which significantly complicate the subsequent research. Existing underwater image enhancement methods often treat the haze and color cast as a unified degradation process and disregard their independence and interdependence, which limits the performance improvement. Here, we propose a Vision Transformer (ViT)-based network (referred to as WaterFormer) to improve the underwater image quality. WaterFormer contains three major components: a dehazing block (DehazeFormer Block) to capture the self-correlated haze features and extract deep-level features, a Color Restoration Block (CRB) to capture self-correlated color cast features, and a Channel Fusion Block (CFB) to capture fusion features within the network. To ensure authenticity, a soft reconstruction layer based on the underwater imaging physics model is included. To improve the quality of the enhanced images, we introduce the Chromatic Consistency Loss and Sobel Color Loss to train the network. Comprehensive experimental results demonstrate that WaterFormer outperforms other state-of-the-art methods in enhancing underwater images.
@@ 1 Introduction
The ocean holds vast natural resources, so deep-sea exploration and development are crucial. Underwater imaging serves as a vital carrier of marine information, and the image quality greatly influences marine engineering. However, with its floating particles, light scattering, and selective light absorption, the deep ocean environment introduces noise, blur, and color distortion [1], which have detrimental effects on underwater detection, underwater visual perception, and other underwater tasks. Underwater image enhancement (UIE) aims to mitigate these issues, restore the true colors, enhance the details, and consequently increase the data reliability and usability. Therefore, research on UIE algorithms is necessary, urgent, and of great significance.
@@
In the early days, specialized hardware [2], multiple different images [3] or special imaging methods [4] were used to enhance the quality of underwater images. Although these methods were effective in certain scenarios, their expensive hardware and complex conditions limited their widespread use. To achieve convenience and universality, increasingly many image processing and computer vision techniques are used to address the challenges of underwater imaging. These techniques can be broadly categorized into non-physical model-based methods, physical model-based methods, and data-driven methods. Non-physical model-based methods do not rely on underwater optical propagation models. They adjust the pixels of the image in the spatial or transform domain to improve the visual quality of the image [5, 6, 7]. Although non-physical model-based methods are convenient and easy to use, they have controversial fidelity due to the lack of physical priors. Moreover, under special conditions, these methods can cause more severe color distortion and artifacts. Physical model-based methods consider the process of light propagation underwater, make assumptions, and build corresponding mathematical models to enhance underwater images. Example physical model-based methods estimate the light scattering parameters [8], medium transmission parameters [9, 10]. Physical model-based methods are image restoration methods and offer higher fidelity. However, they can fail in complex underwater environments and challenging imaging conditions when the assumptions do not hold. In recent years, data-driven techniques have gained significant momentum, and more researchers have begun using deep learning to enhance underwater images [11, 12]. With their large parameter space and powerful learning capabilities, deep learning models often outperform traditional methods when ample datasets are available. However, data-driven methods have three challenges. First, most data-driven methods face a trade-off between performance and the number of parameters. Second, they separately consider features of the underwater environment or rely on a single degradation process to guide the network, which limits their UIE performance. Third, they often cannot guarantee the authenticity of the enhanced data.
@@
Hence, we devised WaterFormer, which is a lightweight but high-performance network that considers the atomization and color cast as inherent primary degradation features to the underwater environment to restore images. Specifically, WaterFormer uses the DehazeFormer Block proposed in  [13] to achieve self-attention for the haze features and extract deep features. Using the channel-wise self-attention matrix transformation, our Color Restoration Block (CRB) performs self-attention on color cast features to help the network infer results with a more realistic color distribution. We also introduced the Channel Fusion Block (CFB), which synthesized the global information from different features and refined it to generate more effective fusion features. Considering the constraints of the underwater imaging physical model, we incorporated an underwater soft reconstruction layer at the end of the network to enhance the fidelity of the results. Furthermore, we introduced the Chromatic Consistency Loss and Sobel Color Loss to train the network in maintaining chromatic consistency, preserving fine color details, and enhancing the image quality and generalization across various datasets.
@@
  • A novel end-to-end network to enhance underwater images is proposed. Our approach uses the DehazeFormer Block, successfully transforms haze features and improves the image deblurring capability. It introduces an underwater soft reconstruction layer to ensure the authenticity of the results.
@@ 2.1 Traditional Methods
Traditional methods can be broadly categorized into methods based on non-physical models and those based on physical models. Non-physical-model-based methods do not involve modeling the underwater optical transmission process. Instead, they adjust the pixels of an image in the spatial domain or transform domain using the characteristic statistics and visual perceptual analysis. These methods include contrast enhancement [14], histogram equalization [15, 16], white balance [17], wavelet transform [18], fusion-based methods [19, 20], retinex-based methods [21, 22, 23], and others. For example, Hitam et al. [16] applied the mixture Contrast Limited Adaptive Histogram Equalization to underwater images in the RGB and HSV color spaces. Singh et al. [18] performed a discrete wavelet transform on images and estimated the color correction factors that were associated with specific color casts to adjust the input image pixels. Ancuti et al. [19] separately conducted color correction and contrast enhancement on underwater images and subsequently fused the two versions based on automatically calculated weights. Although most non-physical model-based methods are simple to implement and computationally efficient, they lack prior knowledge of the optical propagation, which can result in oversaturated or undersaturated color distributions.
Physical-model-based methods are commonly used in underwater image enhancement. They involve mathematically modeling the underwater degradation process, estimating the parameters of the degradation model, and ultimately recovering clear images. These methods address various aspects of underwater image degradation, such as estimating the parameters related to the light scattering [8] and medium transmission [9, 24]. For example, Hou et al. [8] used optical response functions, including Point Spread Function (PSF) and Modulation Transfer Function (MTF), and deconvolution techniques for image restoration. Ke et al. [10] considered the color, saturation, and detail information in images to construct scene depth maps and edge maps to estimate the medium transmission rates. Galdran et al. [25] adjusted the Dark Channel Prior (DCP) method by incorporating red channel priors to estimate the medium transmission map for underwater scenes. Wang et al. [24] introduced the adaptive attenuation-curve prior to estimate medium transmission parameters. Physical-model-based methods are based on prior assumptions and offer high fidelity in image restoration. However, their applicability is often limited to specific underwater degradation phenomena, so they have lower generality.
@@ 2.2 Data-driven Methods
The use of deep learning techniques in UIE tasks has been a major trend in recent years. These approaches mainly follow two paradigms: the encoding-decoding paradigm and the generative paradigm. On one hand, methods based on the encoding-decoding paradigm typically encode images into a latent space and continuously learn low-level features and deep representations of images to reconstruct the final results through a decoder. Encoding-decoding methods often use global residual connections or gate fusion mechanisms to facilitate network learning. Li et al. [26] introduced a gate fusion network called Water-Net, which enhanced images by fusing the original image with white-balanced, histogram-equalized, and gamma-corrected versions. Water-Net can adapt to different underwater environments and enhance the image. Xue et al. [27] analyzed underwater degradation factors from two perspectives: color distortions and veil effects. They developed the Multi-Branch Multi-Variable Network (MBANet) to restore underwater images. Li et al. [28] proposed an enhancement network called Ucolor, which operated in the HSV, RGB, and Lab color spaces to encode images and decodes image features with a medium transmission map. Therefore, Ucolor can learn and selectively enhance different visual representations of images.
On the other hand, Generative Adversarial Networks (GANs) are key components of the generative paradigm, and they mainly consist of a generator and a discriminator. The generator produces samples from random noise, and the discriminator distinguishes between real and generated samples. Through adversarial learning, they iteratively improve each other until the generator can generate realistic samples. Islam et al. [29] designed an adversarial network called FUnIE based on conditional generation, where the generator was constructed using the Unet [30] architecture, and the discriminator was a Markovian discriminator [31]. Cong et al. [32] proposed PUGAN, which guided the generator to produce more interpretable results by estimating medium transmission parameters. Liu et al. [33] split the network training into weak and strong supervision stages, which enhanced its performance while decreasing the dependence on paired training data.
@@
Most studies on UIE tasks are based on Convolutional Neural Networks (CNNs) because they can effectively capture local features in images through convolutional and pooling layers. In recent years, the Visual Transformer (ViT) [34] was proposed, and it can better capture both local and global perspectives of images [35] and gain more similarity between features than CNNs. As a result, some researchers began using ViT to handle UIE tasks. For example, Peng et al. [36] used the ViT architecture to build a generative adversarial paradigm network called U-shape. Shen et al. [37] constructed a ViT network called UDAformer that included a spatial attention mechanism and a channel attention mechanism in an encoder-decoder paradigm.
@@
Although many data-driven algorithms exhibit excellent generalization, they often grapple with a trade-off between parameter count and performance. These methods overlook the interconnected but independent nature of the haze and color distortion phenomena. Treating these features as a singular degradation process neglects the interplay between haze removal and color restoration. Furthermore, existing approaches lack tailored loss functions for underwater image enhancement (UIE), which diminishes the learning capacity of the network. Hence, we developed a lightweight, high-performance network for haze and color distortion features with two proposed loss functions to enhance the underwater image features and optimize the learning capability of the network.
@@ 3 Proposed Method
Fig. 1: Overall structure of WaterFormer, which includes the DehazeFormer Block as the core component for feature extraction and transformation in the network. We enhanced the ability of the network to aggregate local information by incorporating the FReLU activation function in the MLP structure. Additionally, our proposed CRB and CFB contribute to enhancing the color features of underwater images and integrating features from different branches. The entire WaterFormer is a three-stage improved Unet network, where the Down-Sample and Up-Sample stages are implemented using convolution and pixelshuffle, respectively.
@@ 3.1 Overall framework and hierarchical structure
The proposed WaterFormer adopts the U-shaped[30] encoder-decoder architecture as the backbone. In the encoder part, the network uses the DehazeFormer Block structure[13] with self-correlation to refine the haze features and deep features. The Color Restoration Block (CRB) was designed to enhance the color characteristics of the feature maps. In the decoder part, in addition to refined features, we used the CFB for skip connections to efficiently fuse different features at different stages. Finally, to improve the realism of the estimated image, we incorporated an underwater soft reconstruction layer into WaterFormer at the end, which innovated based on the soft reconstruction layer in [13] and replaced the original global residual learning. Specifically, the underwater image formation process in [38] is as follows:
@@
To reduce the complexity and computational cost of the network, we removed a significant number of convolutional layers from the original DehazeFormer Block structure and only retained the convolution operations related to the self-attention and multilayer perceptron (MLP). To compensate for the limited utilization of local information due to the decrease in number of convolutional layers, we replaced the activation function in the MLP with a more locally enhanced FReLU [39] and designed WaterFormer as a three-stage Unet-based network. If there are more complex underwater environments and less demand for network speed performance, the number of encoding-decoding stages can be increased to enhance the network performance.
@@ 3.1.1 Color Restoration Block (CRB)
Due to the absorption and scattering of light in water, underwater images tend to appear bluish or greenish compared to land images, which is a channel-level per-pixel interference. Therefore, leveraging the global information among the channels for color balancing is an important task in underwater image enhancement. Inspired by the matrix transformations in  [40, 41], we designed the CRB to perform matrix transformations on the pixel-wise information among input feature channels using a cross-channel cross-covariance. This process generates an attention map that implicitly encodes the global context, facilitates interactions within the channels and promotes the color enhancement. The computational complexity of the channel-level global self-attention is linear, so the CRB can exhibit faster computational performance than spatial self-attention computations [42, 34, 35]. Fig. 2 illustrates the specific structure of the CRB.
@@
Here, $PConv(\cdot)$ is a $1\times 1$ point-wise convolution; $DConv(\cdot)$ is a $3\times 3$ depth-wise convolution; $Flat(\cdot)$ is Flattening operation; $Q,K,V\in{{\mathbb{R}}^{c\times hw}}$.

### v2

@@ Abstract
Underwater visual imaging is crucial for marine engineering, but it suffers from low contrast, blurriness, and color degradation, which hinders downstream analysis. Existing underwater image enhancement methods often treat the haze and color cast as a unified degradation process, neglecting their inherent independence while overlooking their synergistic relationship. To overcome this limitation, we propose a Vision Transformer (ViT)-based network (referred to as WaterFormer) to improve underwater image quality. WaterFormer contains three major components: a dehazing block (DehazeFormer Block) to capture the self-correlated haze features and extract deep-level features, a Color Restoration Block (CRB) to capture self-correlated color cast features, and a Channel Fusion Block (CFB) that dynamically integrates these decoupled features to achieve comprehensive enhancement. To ensure authenticity, a soft reconstruction layer based on the underwater imaging physics model is included. Further, a Chromatic Consistency Loss and Sobel Color Loss are designed to respectively preserve color fidelity and enhance structural details during network training. Comprehensive experimental results demonstrate that WaterFormer outperforms other state-of-the-art methods in enhancing underwater images.
@@ 1 Introduction
The ocean holds vast natural resources, so deep-sea exploration and development are crucial. Underwater imaging serves as a vital carrier of marine information and image quality greatly influences marine engineering. However, with its floating particles, light scattering and selective light absorption, the deep ocean environment introduces noise, blur, and color distortion [1], which have detrimental effects on underwater detection, underwater visual perception, and other underwater tasks [2, 3]. Underwater Image Enhancement (UIE) aims to mitigate these issues, restore the true colors, enhance the details, and consequently increase the reliability and usability of the data. Therefore, research on UIE algorithms is necessary, urgent, and of great significance.
@@
In the early days, specialized hardware [4], multiple different images [5] or special imaging methods [6] were used to improve the quality of underwater images. Although these methods were effective in certain scenarios, their expensive hardware and complex conditions limited their widespread use. To achieve convenience and universality, increasingly many image processing and computer vision techniques are used to address the challenges of underwater imaging. These techniques can be broadly categorized into non-physical model-based methods, physical model-based methods, and data-driven methods. Non-physical model-based methods do not rely on underwater optical propagation models. They adjust in the spatial or transform domain to improve the visual quality of the image [7, 8]. Non-physical model-based methods are user-friendly but may exacerbate color distortion or introduce unnatural artifacts due to their reliance on statistical/empirical adjustments, which ignore the physics of wavelength-selective light attenuation underwater. Moreover, most approaches decouple dehazing and color correction into separate stages, causing iterative error accumulation: information lost in earlier stages cannot be retrieved later. Physical model-based methods enhance underwater images by analyzing the principles of light propagation, formulating assumptions, and constructing mathematical models to simulate underwater processes. For example, these approaches often estimate critical parameters such as light scattering coefficients [9] and medium transmission rates [10, 11]. Using the analogous modeling frameworks shared between underwater imaging and atmospheric haze imaging, these methods achieve simultaneous color restoration and dehazing with high fidelity. However, their effectiveness diminishes in complex underwater environments due to the overdependence of simplified assumptions, such as uniform illumination and static water conditions, which fail to account for real-world challenges such as artificial directional lighting, dynamic turbulence, and suspended particulate matter. Additionally, the interdependent nature of color correction and dehazing introduces mutual performance degradation when ideal conditions for either task (e.g., accurate parameter estimation or stable scene properties) are not met. In recent years, data-driven techniques have gained significant momentum and more researchers have begun to use deep learning to enhance underwater images [12, 13]. With their large parameter space and powerful learning capabilities, deep learning models often outperform traditional methods when ample datasets are available. However, data-driven methods face three challenges. First, most data-driven methods face a trade-off between performance and the number of parameters. Second, most of them only focus on deep feature extraction or color-based guidance, neglecting haze features’ importance, which limits their UIE performance in dehazing and edge detail retention. Third, they often cannot guarantee the authenticity of the enhanced data due to the lack of physical model guidance.
@@
Hence, we devised WaterFormer, which is a lightweight but high-performance network that considers the haze and color cast as inherent primary degradation features of the underwater environment to restore images. Specifically, WaterFormer uses the DehazeFormer Block proposed in [14] to achieve self-attention for the haze features and extract deep features. Using the channel-wise self-attention matrix transformation, our Color Restoration Block (CRB) performs self-attention on color cast features to help the network infer results with a more realistic color distribution. We also introduced the Channel Fusion Block (CFB), which synthesized the global information from different features and refined it to generate more effective fusion features. Considering the constraints of the underwater imaging physical model, we incorporated an underwater soft reconstruction layer at the end of the network to enhance the fidelity of the results. Furthermore, we introduced the Chromatic Consistency Loss and Sobel Color Loss to train the network to maintain chromatic consistency, preserve fine color details, and enhance image quality and generalization across various datasets.
@@
  • A novel end-to-end network is proposed to enhance underwater images. Our approach uses the DehazeFormer Block, successfully transforms haze features, and improves the image deblurring capability. It introduces an underwater soft reconstruction layer to ensure the authenticity of the results.
@@ 2.1 Traditional Methods
Traditional methods can be broadly categorized into methods based on non-physical models and those based on physical models. Non-physical model-based methods do not involve modeling the underwater optical transmission process. Instead, they adjust the pixels of an image in the spatial domain or transform domain using the characteristic statistics and visual perceptual analysis. These methods include contrast enhancement [15], histogram equalization [16], wavelet transform [17], fusion-based methods [18], retinex-based methods [19], and others. For example, Hitam et al. [16] applied the mixture Contrast Limited Adaptive Histogram Equalization to underwater images in the RGB and HSV color spaces. Singh et al. [17] performed a discrete wavelet transform on images and estimated the color correction factors that were associated with specific color casts to adjust the input image pixels. Ancuti et al. [18] separately conducted color correction and contrast enhancement on underwater images and subsequently fused the two versions based on automatically calculated weights. Although most non-physical model-based methods are simple to implement and computationally efficient, they lack prior knowledge of the optical propagation, which can result in oversaturated or undersaturated color distributions.
Physical model-based methods are commonly used in UIE. They involve mathematical modeling of the underwater degradation process, estimating the parameters of the degradation model, and ultimately recovering clear images. These methods address various aspects of underwater image degradation, such as estimating the parameters related to the light scattering [9] and medium transmission [10, 20]. For example, Hou et al. [9] used optical response functions, including Point Spread Function (PSF) and Modulation Transfer Function (MTF), and deconvolution techniques for image restoration. Ke et al. [11] considered the color, saturation, and detail information in images to construct scene depth maps and edge maps to estimate the medium transmission rates. Galdran et al. [21] adjusted the Dark Channel Prior (DCP) method by incorporating red channel priors to estimate the medium transmission map for underwater scenes. Physical model-based methods are based on prior assumptions and offer high fidelity in image restoration. However, their applicability is often limited to specific underwater degradation phenomena, so they have lower generality.
@@ 2.2 Data-driven Methods
The use of deep learning techniques in UIE tasks has been a major trend in recent years. These approaches mainly follow two paradigms: the encoding-decoding paradigm and the generative paradigm. On the one hand, methods based on the encoding-decoding paradigm typically encode images into a latent space and continuously learn low-level features and deep representations of images to reconstruct the final results through a decoder. Encoding-decoding methods often use global residual connections or gate fusion mechanisms to facilitate network learning. Li et al. [13] introduced a gate fusion network called Water-Net, which enhanced images by fusing the original image with white-balanced, histogram-equalized, and gamma-corrected versions. Water-Net can adapt to different underwater environments and enhance the image. Xue et al. [22] analyzed underwater degradation factors from two perspectives: color distortions and veil effects. They developed the Multi-Branch Multi-Variable Network (MBANet) to restore underwater images. Li et al. [23] proposed an enhancement network called Ucolor, which operated in the HSV, RGB, and Lab color spaces to encode images and decode image features with a medium transmission map. Therefore, Ucolor can learn and selectively enhance different visual representations of images.
On the other hand, Generative Adversarial Networks (GANs) and Diffusion Models constitute two pivotal methodologies within the generative paradigm. GANs comprise two components: a generator that synthesizes samples from random noise, and a discriminator that differentiates authentic samples from synthetic ones. Through adversarial optimization, both components co-evolve until the generator attains photorealistic synthesis capability. Diffusion models involve a forward diffusion process and a reverse generation stage. During the forward process, incremental Gaussian noise is systematically introduced to an image until its statistical properties converge to a standard normal distribution, concurrently training a Unet [24] to estimate the noise profile. The reverse phase leverages the trained Unet to progressively refine the corrupted latent representation into a coherent image through iterative denoising operations. Islam et al. [25] designed an adversarial network called FUnIE based on conditional generation, where the generator was constructed using Unet architecture and the discriminator was a Markovian discriminator. Cong et al. [26] proposed PUGAN, which guided the generator to produce more interpretable results by estimating medium transmission parameters. Tang et al. [27] developed a diffusion model-based UIE framework, employing piecewise sampling and evolutionary algorithm search to optimize generation efficiency. Zhao et al. [28] innovatively integrated frequency domain representations into the diffusion architecture, where a Wavelet-based Fourier interaction network and Frequency Residual Diffusion Adjustment Module collectively demonstrated the critical role of spectral characteristics in UIE tasks.
@@
Most studies on UIE tasks are based on Convolutional Neural Networks (CNNs) because they can effectively capture local features in images through convolutional and pooling layers. In recent years, the ViT [29] was proposed, and it can better capture both local and global perspectives of images and gain more similarity between features than CNNs. As a result, some researchers began using ViT to handle UIE tasks. For example, Peng et al. [30] used the ViT architecture to build a generative adversarial paradigm network called U-shape. Shen et al. [31] constructed a ViT network called UDAformer that included a spatial attention mechanism and a channel attention mechanism in an encoder-decoder paradigm.
@@
Although many data-driven algorithms exhibit excellent generalization, they often grapple with a trade-off between parameter count and performance. These methods overlook the interconnected but independent nature of the haze and color distortion phenomena. Treating these features as a singular degradation process neglects the interplay between haze removal and color restoration. Furthermore, existing approaches lack tailored loss functions for UIE, which decreases the learning capacity of the network. Hence, we developed a lightweight, high-performance network for haze and color distortion features with two proposed loss functions to enhance the underwater image features and optimize the learning capability of the network.
@@ 3 Proposed Method
Fig. 1: Overall structure of WaterFormer, which includes the DehazeFormer Block as the core component for feature extraction and transformation in the network. It’s ability to aggregate local information is enhanced through the incorporation of the FReLU in the MLP. Additionally, the proposed CRB and CFB are employed to improve the color features of underwater images and facilitate the integration of features across different branches. WaterFormer is built as a three-stage, enhanced Unet network, with Down-Sample and Up-Sample stages implemented using convolution and pixelshuffle, respectively. Input/output images are from the EUVP dataset[25].
Block Type Num. of Blocks Embedding Dims Num. of Heads MLP Ratio Act. Function of MLP
DehazeFormer Block [2, 4, 2] [24, 48, 24] [3, 6, 3] [3, 4, 3] FReLU
Color Restoration Block [2, 2] [24, 48] [3, 6] [3, 4] FReLU
Table 1: Detailed structure of the Transformer block.
@@ 3.1 Overall framework and hierarchical structure
The proposed WaterFormer adopts the U-shaped[24] encoder-decoder architecture as the backbone. In the encoder part, the network uses the DehazeFormer Block[14] with self-correlation to refine the haze features and deep features. The Color Restoration Block (CRB) was designed to enhance the color characteristics of the feature maps. In the decoder part, in addition to refined features, we used the CFB for skip connections to efficiently fuse different features at different stages. Finally, to improve the realism of the estimated image, we incorporated an underwater soft reconstruction layer into WaterFormer at the end, which innovated based on the soft reconstruction layer in [14] and replaced the original global residual learning. Specifically, the underwater image formation process in [32] is as follows:
@@
To reduce the complexity and computational cost of the network, we removed a significant number of convolutional layers from the original DehazeFormer Block structure and only retained the convolution operations related to the self-attention and multilayer perceptron (MLP). To compensate for the limited utilization of local information due to the decrease in number of convolutional layers, we replaced the activation function in the MLP with a more locally enhanced FReLU [33] and designed WaterFormer as a three-stage Unet-based network. If there are more complex underwater environments and less demand for network speed performance, the number of encoding-decoding stages can be increased to enhance the network performance. Fig. 1 and Tab. 1 present the detailed structure of WaterFormer.
@@ 3.1.1 Color Restoration Block (CRB)
Due to the absorption and scattering of light in water, underwater images tend to appear bluish or greenish compared to land images, which is a channel-level per-pixel interference. Therefore, leveraging the global information among the channels for color balancing is an important task in UIE. Inspired by the matrix transformations in  [34, 35], we designed the CRB to perform matrix transformations on the pixel-wise information among input feature channels using a cross-channel cross-covariance. This process generates an attention map that implicitly encodes the global context, facilitates interactions within the channels and promotes the color enhancement. The computational complexity of the channel-level global self-attention is linear, so the CRB can exhibit faster computational performance than spatial self-attention computations [29]. Fig. 2 illustrates the specific structure of the CRB.
@@
Here, $PConv(\cdot)$ is a $1\times 1$ point-wise convolution; $DConv(\cdot)$ is a $3\times 3$ depth-wise convolution; $Flat(\cdot)$ is a flattening operation; $Q,K,V\in{{\mathbb{R}}^{c\times hw}}$.